OpenStudy (anonymous):

Given triangle GHI with G(4.-3), H(-4,2), and I(2,4), find the perpendicular bisector of line HI in standard form. I really need help on this.....I'm totally confused.

OpenStudy (anonymous):



OpenStudy (anonymous):

first find the slope of the of the line HI, you can use the point-slope formula to do this: \[y_2-y_1 = m*(x_2-x_1)\] lets rename m to m_a once you found the slope, m_a, for the line HI, find the slope of a line that would be perpendicular to it. the perpendicular slope we will call m_b. use the equation for perpendicular slopes: \[m_a * m_b = -1\] plug in m_a, and solve the equation for m_b. you now have the slope of the line that is perpendicular to the line HI. Take m_b, and point 'G'. plug those values into the standard slope-intercept form equation of y = mx+b solve for b. once you have 'b' re write the equation of y=mx+b, but only plug in the value for 'm' and the value for 'b'

OpenStudy (anonymous):

Oh my gosh! thank you so much!

OpenStudy (anonymous):

Hold on.. I'm stuck on the last part. When i first write y=mx+b what do I plug in where??

OpenStudy (anonymous):

y=mx+b for x and y, plug in the point G, so (4, -3) for m, plug in what you solved for as the slope of the perpendicular line slope (sorry, OpenStudy isn't showing notifications today)

OpenStudy (anonymous):

aha, thank you!!

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

...I got y=-1 1/3x+7/3 haha, did i go wrong somewhere?

OpenStudy (anonymous):

ya, there was a little mix up somewhere, so the slope of the HI line is m = 1/3 or m_a = 1/3 then m_b = -3 do you see where u slipped up?

OpenStudy (anonymous):

Yeah! i got mixed up in the second equation... Ok, i got it now, y=-3x+9??

OpenStudy (anonymous):

excellent work, so right now y=-3x+9 is in slope intercept form so re write the equattion into standard form Ax + By = C

OpenStudy (anonymous):

Whoa, what.

OpenStudy (anonymous):

haha, so just rearrange the equation y=-3x+9 so that 9 is by itself

OpenStudy (anonymous):

i'm logging off, hope you got it figured out!

OpenStudy (anonymous):

thank you (: I got it.